x+y+z=6
2x-y+z=3
3x-z=0

Answer by tommyt3rd(5050)

from 3:
z=3x
1 plus 2:
(2x-y+z)+(x+y+z)=3+6
3x+2z=9
3x+2(3x)=9
9x=9
x=1
so z=3
from 1:
x+y+z=1+y+3=6
y+4=6
y=2

Solution: (1,2,3)


:)